OpenStudy (anonymous):

Mrs. Wang is 23 years older than her daughter. In 5 years, their age will total 63. How old are they now?

OpenStudy (lgbasallote):

let x = daughter's age x + 23 = mrs. wang's age in 5 years daughter's age would be x + 5 in 5 years mrs. wang's age would be x + 23 + 5 the sum of those is 63 so... x + 5 + x + 23 + 5 = 63 2x + 33 = 63 2x = 63 - 33 2x = 30 can you solve for x?

OpenStudy (anonymous):

thank you so much

OpenStudy (lgbasallote):

<tips hat>